OpenStudy (anonymous):

Logic/math? I don't understand what its asking! If 12 = 10 and 32 = 26, then 22 =

OpenStudy (anonymous):

The left side is in base 8, the right side is in base 10. So instead of the place values being 1, 10, 100, etc., the place values are 1, 8, 64, etc. In the first example, 12 = (1*8) + (2*1) = 10, and then 32=(3*8)+(2*1)=26, so 22=(2*8)+(2*1)=18.

OpenStudy (anonymous):

Oh, ok:) I was kind of confused at first but this helps a lot! Thanks!